The physician’s order reads: Aminophylline 80 mg/hr IV. The drug is supplied in a bag containing 500 mg of Aminophylline in 50 mL of solution.

a) What is the concentration (mg/mL) of Aminophylline in the bag?

b) At what rate (mL/hr) would you deliver the medication?

10 mg/mL , 8 mL/hr
50mg/ml, 30ml/hr
50mg/ml, 10ml/hr
10mg/ml, 10ml/hr
Answer #1

Given:

Medication :Aminophylline

Route=IV

Volume :50ml

Available :500mg in 50ml

Order=80mg/hr

a)when 500mg=50ml

Then ??=1ml

Cross multiply and divide

=500×1÷50

=10mg

The concentration is 10mg/ml of Aminophylline

b)Flow rate (ml/hr)=Volume(ml) ÷Time(hr)

When 80mg=1hour

Then 500mg=??

=1×500÷80

=6.25 or 6gtt /min

Now flow rate =Volume (ml)÷Time(hr)

6=50÷T

T=50÷8

T=8.3 or 8ml /hr

Ans.10mg/5ml,8ml/hr